All structures in this assignment are trusses (all assumptions about trusses hold)

Use method of sections to obtain answers. Cutting right next to a joint doesn't count.

Reactions are given as horizontals "H" and verticals "V": positive = up/right; negative = down/left

5. Find forces in members c-e and f-d.

5. Find forces in members c-e and f-d. Note that H = -30k, Vi= -30k, Vh= +50k 1104104 51110* 15k 10 30 | 15 15
Answer #1

skip, lokip skop lokip to by cho T 2 Fx = 0 = f = 20-410 = 30 kip Ify = 0 7 futhu = 5+10+5 = 20 kip Im=0 fu (30)+10(60) +20 (